One card is randomly selected from a deck of cards. Find the odds against drawing a

red

seven

.


1
Expert's answer
2020-11-12T18:15:41-0500

Answer 1)


Odds against drawing a red card = "\frac {Probability\ of\ drawing\ non-red\ card} {Probability\ of\ drawing\ red\ card}"


There are total 26 red cards & 26 non-red cards in a pack of 52 cards.


Probability of drawing a non-red card = "\frac {26C1}{52C1} = \frac {26}{52}=0.5"


Probability of drawing a red card = "\frac {26C1}{52C1} = \frac {26}{52}=0.5"


Hence the odds against drawing a red card = "\frac {0.5}{0.5}=1:1"


Answer 2)


Odds against drawing a seven = "\frac {probability\ of\ not\ drawing\ seven }{probability\ of\ drawing\ seven }"


There are total 4 Seven cards in a pack


Probability of not drawing a seven = "\frac {48C1}{52C1}=\frac {48}{52}"


Probability of drawing a seven = "\frac {4C1}{52C1}=\frac {4}{52}"


Hence the odds against drawing a seven = "\frac {\frac {48}{52}}{\frac {4}{52}}=\frac {48*52}{4*52}=\frac {48}{4}=12:1"